[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Apr 28, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Apr 28 15:07:56 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt Sardi

----------

TAFB:
1. NH2/Atlantic 18z NWPS run only came into GFE through 60 hours yesterday.  
Chris Mello is investigating.
HSU:
1. TSB identified and NCO fixed issues with TILEFISH X server config file  
(xorg.conf), which could have caused N-AWIPS performance degradation.
Other items:
1. NCO is scheduled to update to NHC's iptables on the 112 VLAN today at  
1600Z. This change should not affect NHC networking.

2. There has been a recurring problem with the LHCP channel of NHC's GOES-E  
rack, which ISatSS pulls GOES 16 data from.

Harris will be on-site starting Tuesday 5/2 to try to resolve the issue;  
until then, NHC will be pulling GOES16 data from the GOES-Spare rack.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170428/16b68723/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list